# Extracting work from coherence in a two-mode Bose-Einstein condensate

### v1

@@ Abstract
We show how work can be extracted from number-state coherence in a two-mode Bose-Einstein condensate. With careful tuning of parameters, a sequence of thermodynamically reversible steps transforms a Glauber coherent state into a thermal state with the same energy probability distribution. The work extracted during this process arises entirely from the removal of quantum coherence. More generally, we characterise quantum (from coherence) and classical (remaining) contributions to work output, and find that in this system the quantum contribution can be dominant over a broad range of parameters. The proportion of quantum work output can be further enhanced by squeezing the initial coherent state. Due to the many-body nature of the system, the work from coherence can equivalently be understood as work from entanglement.
@@ I Introduction
Building on the generic protocol in [11], in this manuscript we show how number-state coherence in a two-mode BEC can empower work extraction. We introduce a formalism to quantify the work directly available from number-state coherence and the work that would be available classically. To obtain work that is predominantly quantum in origin, the initial state must be close to thermal when projected onto the energy eigenbasis. This can be achieved by using a Glauber coherent state with mean boson number less than a critical value that depends on temperature. We show how squeezing the initial state can substantially increase the proportion of quantum work extracted. The number-state coherence in our protocol arises from entangling correlations between the two modes of the BEC, and hence the work output can equivalently be interpreted as work from entanglement. To our knowledge, our work presents the first proposal showing how work can be extracted from coherence in a two-mode BEC, paving the way for experimental demonstration and highlighting the utility of two-mode BECs in thermodynamic applications.
@@ II.1 Extracting work from coherence
$W=\beta^{-1}(S^{E}-S^{V})=\beta^{-1}D(\rho||\rho_{\mathrm{ed}}).$ (4)
@@ II.2 System setup
We consider a two-mode BEC with fixed total boson number $N$. The two modes could be hyperfine spin states or spatial modes. Number states of the system are denoted by |$\ket{n,k}$, with $n$ and $k=N-n$ the number of bosons in the two modes respectively. We denote the annihilation operators for the two modes by $\hat{a}$ and $\hat{b}$, with $\hat{a}\ket{n,k}=\sqrt{n}\ket{n-1,k}$ and $\hat{b}\ket{n,k}=\sqrt{k}\ket{n,k-1}$. The Hamiltonian for the system is ($\hbar\equiv 1$ here and throughout) [35]
$\begin{split}\hat{H}=&\omega_{1}\hat{a}^{\dagger}\hat{a}+\omega_{2}\hat{b}^{\dagger}\hat{b}+g_{11}\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}+g_{22}\hat{b}^{\dagger}\hat{b}^{\dagger}\hat{b}\hat{b}\\ &+g_{12}\hat{a}^{\dagger}\hat{b}^{\dagger}\hat{a}\hat{b}+E_{0}\\ \equiv&\Delta\hat{a}^{\dagger}\hat{a}+g\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}.\end{split}$ (7)
Here $\omega_{i}$ is the single-particle energy of mode $i$ and $g_{ij}$ is the interaction energy between particles from modes $i,j$, with $i,j\in\{1,2\}$. The vacuum energy is $E_{0}$. The operator $\hat{N}=\hat{a}^{\dagger}\hat{a}+\hat{b}^{\dagger}\hat{b}$ commutes with $\hat{H}$ and hence the total boson number is conserved. In the second line of Eq. (7) we replace $\hat{b}^{\dagger}\hat{b}$ by $\hat{N}-\hat{a}^{\dagger}\hat{a}$ and define $\Delta\equiv\omega_{1}-\omega_{2}-2(N-1)g_{22}+(N-1)g_{12}$ and $g\equiv g_{11}+g_{22}-g_{12}$. We choose $E_{0}=-N\omega_{2}-N(N-1)g_{22}$ so that $\hat{H}\ket{0}=0$ for convenience. The choice of $E_{0}$ will not affect the net output of any cyclic process, for example the performance of an engine cycle. The detuning $\Delta$ for two spatial modes can be controlled via tuning the trapping potential [46], and for two spin states via tuning of a Zeeman field [47, 48]. As will be seen below, we will require immiscible interactions $g>0$ [49]. Interactions between the two modes (strength $g_{12}$) are included for generality, but are not necessary.
@@
the partition function. Due to particle interactions, the boson number distribution of a thermal state is approximately Gaussian (when $-\Delta\gg g>0$), in contrast to the Planck distribution of non-interacting bosons. This feature will be essential for the efficacy of our protocol as it will enable close matching between the initial and thermal boson number distributions. We assume $\beta g\ll 1$; noting that $gN\sim\mu$, where $\mu$ is the BEC chemical potential, this is ensured when $\beta\mu/N\ll 1$, a condition satisfied in dilute gaseous BECs [50, 51]. We can then approximate the sum in Eq. (10) by an integral. We also assume $N\gg 1$ and $\langle\hat{a}^{\dagger}\hat{a}\rangle\ll N$, i.e. most of the particles are in the $\hat{b}$ mode throughout the protocol. We can then set $N\rightarrow\infty$ and approximate |$\ket{n,N-n}$ by a harmonic oscillator number state, which simplifies our calculations [52, 53]. Evaluating Eq. (10) with these two approximations gives
$Z\approx\sqrt{\frac{\pi}{\beta g}}e^{\beta(g-\Delta)^{2}/4g}\left(\frac{1+\operatorname{erf}\left(\frac{\beta(g-\Delta)}{2\sqrt{\beta g}}\right)}{2}\right).$ (11)
@@ III Results
$p_{n}^{G}=\frac{|\alpha|^{2n}e^{-|\alpha|^{2}}}{n!}\approx\frac{1}{\sqrt{2\pi|\alpha|^{2}}}e^{-(n-|\alpha|^{2})^{2}/2|\alpha|^{2}}.$ (13)
@@
For more general parameters, the work output will have both a contribution from coherence and a “classical” contribution. Below we present an explicit protocol to show how work can be extracted from the coherence of a GCS. We then show how the work output depends on particle number and temperature, identifying a regime where the work output is dominated by the contribution from coherence. Next we show how the work output can be improved by squeezing the initial GCS and finally we discuss the role of entanglement in the system.
@@ III.1 Protocol to extract work from a GCS
with $S_{i}^{V}(S_{f}^{V})$ the initial (final) von Neumann entropies of the system. We decompose $W$ into two parts: A classical part $W_{\mathrm{class}}$ and a quantum part $W_{\mathrm{quant}}$,
$\begin{split}W_{\mathrm{class}}=&U_{i}-U_{f}-\beta^{-1}(S^{E}_{i}-S^{E}_{f}),\\ W_{\mathrm{quant}}=&\beta^{-1}(S^{E}_{i}-S^{V}_{i}).\end{split}$ (16)
Here $S^{E}_{i}(S^{E}_{f})$ is the initial (final) energy entropy, see Eq. (2). Equation (16) decomposes the total entropic change $S^{V}_{f}-S^{V}_{i}$ into a classical contribution $S^{E}_{f}-S^{E}_{i}$ and a quantum contribution $S^{E}_{i}-S^{V}_{i}=D(\rho||\rho_{\mathrm{ed}})$ arising from coherence, Eq. (3). For the protocol in Fig. 1 we have $S^{V}_{i}=0$ since the initial state is pure and $S^{E}_{f}=S^{V}_{f}$ since the final state is thermal. Note that a projection of the initial state Eq. (12) onto the energy eigenbasis prior to work extraction would result in $W_{\mathrm{quant}}=0$, but would not change $W_{\mathrm{class}}$. The protocol in Fig. 1 avoids this measurement, allowing $W_{\mathrm{quant}}$ to be extracted. A decomposition similar to Eq. (16) was applied to ergotropy in [13].
@@ III.2 Classical and quantum contributions to the work extracted from a GCS
With an initial state (12), the outputs $W_{\mathrm{quant}}$ and $W_{\mathrm{class}}$ in Eq. (16) can be calculated analytically. We choose $|\alpha|$ so that the GCS has mean boson number equal to the thermal state, $|\alpha|^{2}=\langle\hat{n}\rangle_{\mathrm{therm}}\equiv\langle\hat{n}\rangle$. Here $\hat{n}=\hat{a}^{\dagger}\hat{a}$ and a “therm” subscript denotes expectation values for state $\rho_{\mathrm{therm}}$, Eq. (8). We have checked that this choice of $|\alpha|$ minimises $W_{\mathrm{class}}$ to a very good approximation over the parameters explored. The energetic and entropic terms in Eq. (16) are
$\begin{split}U_{i}-U_{f}=&-g\left(\langle\delta\hat{n}^{2}\rangle_{\mathrm{therm}}-\langle\hat{n}\rangle\right),\\ S^{E}_{i}=&\frac{1}{2}+\frac{1}{2}\ln 2\pi\langle\hat{n}\rangle+O\left(\frac{1}{\langle\hat{n}\rangle}\right),\\ S^{E}_{f}=&\frac{1}{2}+\frac{1}{2}\ln 2\pi\langle\delta\hat{n}^{2}\rangle_{\mathrm{therm}}+S_{\mathrm{NG}},\end{split}$ (17)
@@
with $u=\frac{\beta(g-\Delta_{0})}{2\sqrt{\beta g}}$ and $G(u)=\frac{e^{-u^{2}}}{\sqrt{\pi}(1+\operatorname{erf}u)}=O(e^{-\beta\Delta_{0}^{2}/g})$. The entropic correction $S_{\mathrm{NG}}$ arises since the thermal distribution is not a perfect Gaussian, and is only important for $u\lesssim 1$.
@@
Again, the $G(u)$ terms are non-Gaussian corrections, which are only important for $u\lesssim 1$. Note for $u\gg 1$ we have $u^{2}\approx\langle\hat{n}\rangle_{\mathrm{therm}}^{2}/2\langle\delta\hat{n}^{2}\rangle_{\mathrm{therm}}$ and hence $u^{-1}$ is essentially the normalized second-order correlation function of the thermal bosonic field [74]. The thermal mean, variance and skewness $\langle\delta\hat{n}^{3}\rangle=\langle(\hat{n}-\langle\hat{n}\rangle\rangle$ are shown in the inset to Fig. 2(a). We focus on cases $\langle\hat{n}\rangle_{\mathrm{therm}}\gtrsim 1$, which is ensured when the initial detuning is less than the reservoir temperature,
@@
Figure 2: (a) Work extracted from a two-mode BEC undergoing the protocol in Fig. 1 for different mean bosonic numbers $\langle\hat{n}\rangle$. Work extracted for an initial GCS (blue curves) divided into quantum $\beta W_{\mathrm{quant}}$ (blue solid curve) and classical $\beta W_{\mathrm{class}}$ (blue dotted curve) contributions, see Eq. (16). The work output is dominated by $W_{\mathrm{quant}}$ for $\langle\hat{n}\rangle\lesssim n_{G}$, with $n_{G}$ given by Eq. (24) and marked by a vertical dashed line. The number state distributions for the GCS and thermal states coincide almost perfectly when $\langle\hat{n}\rangle=1/2\beta g$, at which point $W_{\mathrm{class}}\approx 0$ is minimised (dip in blue dotted curve). Inset: First three cumulants of the thermal boson number distribution, with $\langle\delta\hat{n}^{k}\rangle=\langle(\hat{n}-\langle\hat{n}\rangle)^{k}\rangle$ for $k=2,3$. The distribution is Gaussian for $u\gg 1$, in which case $\langle\hat{n}\rangle_{\mathrm{therm}}\approx u/\sqrt{\beta g}$, $\langle\delta\hat{n}^{2}\rangle_{\mathrm{therm}}\approx 1/2\beta g$ and higher order cumulants are zero. (b) (i)-(iv) Number state distributions for different mean bosonic numbers (four panel values are marked by crosses in (a)) for a GCS (blue curve) and a thermal state (black dashed curve). The GCS distribution is narrower (wider) than the thermal distribution for $\langle\hat{n}\rangle$ smaller (larger) than $\langle\delta\hat{n}^{2}\rangle_{\mathrm{therm}}\approx 1/2\beta g\approx 501$, which results in non-zero $W_{\mathrm{class}}$. Pale red curves in (a) and (b): as for the GCS but for a squeezed coherent state. These are included here for comparison and will be discussed in Sec. III.3. The GCS and squeezed distributions coincide in (ii). All results are for $\beta g=10^{-3}$.
Analytic expressions for $W_{\mathrm{class}}$ and $W_{\mathrm{quant}}$ follow from Eq. (16)–(20) and are plotted in Fig. 2(a). The minimum of the classical work occurs when $p_{n}^{G}$ and $p_{n}^{\mathrm{therm}}$ are as close as possible, see Fig. 2(b). For large $\langle\hat{n}\rangle$ both distributions are approximately Gaussian and hence will be equal when their respective means and variances are equal. Since $p_{n}^{G}$ has equal mean and variance, this requires
@@
This occurs when $\beta\Delta_{0}\approx-1$, at which point $\langle\delta\hat{n}^{2}\rangle_{\mathrm{therm}}\approx 1/2\beta g$. At this optimal point, $W_{\mathrm{class}}=O(\sqrt{\langle\hat{n}\rangle}e^{-\langle\hat{n}\rangle})$ and the work output almost exclusively consists of $W_{\mathrm{quant}}$. (Note our choice of units: plotting work in units of $\beta^{-1}$ is equivalent to plotting the information lost if the work was dissipated as heat, according to Landauer’s principle [75].)
@@
$\begin{split}W_{\mathrm{class}}&\approx g\langle\hat{n}\rangle-\frac{1}{2\beta}\left[1+\ln\left(2\beta g\langle\hat{n}\rangle\right)\right],\\ W_{\mathrm{quant}}&\approx\frac{1}{2\beta}\left[1+\ln\left(2\pi\langle\hat{n}\rangle\right)\right].\end{split}\hskip 28.45274pt(u\gg 1)$ (23)
@@
see Fig. 2(a). Here $\Omega_{k}$ is the $k$th branch of the omega function  [^1] Also known as the Lambert $W$-function, denoted by $W_{k}$. We avoid this notation to avoid confusion with work output.. The function $-\Omega_{-1}(-x)$ satisfies $1\leq-\Omega_{-1}(-x)<\infty$ for $x\in(0,\exp(-1)]$ and monotonically decreases with increasing $x$, with $-\Omega_{-1}(-x)\sim\ln\left(\frac{1}{x}\ln\frac{1}{x}\right)$ for small $x$ [78]. For $\langle\hat{n}\rangle>n_{G}$, the classical work output grows extensively due to the extensive growth of $U_{i}-U_{f}$, see Eq. (23). The quantum work also increases with $\langle\hat{n}\rangle$ due to increased coherence in the initial state. The increase, however, is subextensive, with $W_{\mathrm{quant}}\sim\frac{1}{2\beta}\ln\langle\hat{n}\rangle$. For $\langle\hat{n}\rangle<\langle\delta\hat{n}^{2}\rangle_{\mathrm{therm}}$ (thermal bunching), the number fluctuations of the initial state are less than the final state, see Fig. 2(b). Although this increases $W_{\mathrm{class}}$, we still have $W_{\mathrm{quant}}>W_{\mathrm{class}}$ as long as $\langle\hat{n}\rangle\gg 1$, see Fig. 2(a).

### v3

@@ Abstract
We show how work can be extracted from number-state coherence in a two-mode Bose-Einstein condensate. With careful tuning of parameters, a sequence of thermodynamically reversible steps transforms a Glauber coherent state into a thermal state with the same energy probability distribution. The work extracted during this process arises entirely from the removal of quantum coherence. More generally, we characterise quantum (from coherence) and classical (remaining) contributions to work output, and find that in this system the quantum contribution can be dominant over a broad range of parameters. The proportion of quantum work output can be further enhanced by squeezing the initial state. Due to the many-body nature of the system, the work from coherence can equivalently be understood as work from entanglement.
@@ I Introduction
In this manuscript we show how number-state coherence in a two-mode BEC can empower work extraction. To obtain work that is predominantly quantum in origin the initial state must be close to thermal when projected onto the energy eigenbasis. This is challenging in a quantum many-body system due to the exponentially large number of populations and coherences. We show that this challenge can be overcome by using a Glauber coherent state with mean boson number less than a critical value that depends on temperature. We introduce a formalism to quantify the work directly available from coherence and the work that would be available classically, and show that squeezing the initial state can substantially increase the proportion of quantum work extracted. The number-state coherence in our protocol arises from entangling correlations between the two modes of the BEC. Therefore the work extracted from coherence can equivalently be interpreted as work from entanglement, thus demonstrating the many-body nature of the protocol. To our knowledge, our work presents the first proposal showing how work can be extracted from coherence in a two-mode BEC, paving the way for experimental demonstration and highlighting the utility of two-mode BECs in thermodynamic applications.
@@ II.1 Extracting work from coherence
$W_{\mathrm{quant}}=\beta^{-1}(S^{E}-S^{V})=\beta^{-1}D(\rho||\rho_{\mathrm{ed}}).$ (4)
@@ II.2 System setup
We consider a two-mode BEC with fixed total boson number $N$. The two modes could be hyperfine spin states or spatial modes [46]. Number states of the system are denoted by |$\ket{n,k}$, with $n$ and $k=N-n$ the number of bosons in the two modes respectively. We denote the annihilation operators for the two modes by $\hat{a}$ and $\hat{b}$, with $\hat{a}\ket{n,k}=\sqrt{n}\ket{n-1,k}$ and $\hat{b}\ket{n,k}=\sqrt{k}\ket{n,k-1}$. The Hamiltonian for the system is ($\hbar\equiv 1$ here and throughout) [35]
$\begin{split}\hat{H}=&\omega_{1}\hat{a}^{\dagger}\hat{a}+\omega_{2}\hat{b}^{\dagger}\hat{b}+\frac{g_{11}}{2}\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}+\frac{g_{22}}{2}\hat{b}^{\dagger}\hat{b}^{\dagger}\hat{b}\hat{b}\\ &+g_{12}\hat{a}^{\dagger}\hat{b}^{\dagger}\hat{a}\hat{b}+E_{0}\\ \equiv&\Delta\hat{a}^{\dagger}\hat{a}+g\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}.\end{split}$ (7)
Here $\omega_{i}$ is the single-particle energy and $g_{ii}$ the interaction strength for modes $i=1,2$. We have also allowed for interactions between the two modes of strength $g_{12}$, for example if the two modes are spin states. The vacuum energy is $E_{0}$. The operator $\hat{N}=\hat{a}^{\dagger}\hat{a}+\hat{b}^{\dagger}\hat{b}$ commutes with $\hat{H}$ and hence the total boson number is conserved. In the second line of Eq. (7) we replace $\hat{b}^{\dagger}\hat{b}$ by $\hat{N}-\hat{a}^{\dagger}\hat{a}$ and define
$\Delta\equiv\omega_{1}-\omega_{2}-(N-1)(g_{22}-g_{12})$ (8)
and
$g\equiv\frac{g_{11}+g_{22}-2g_{12}}{2}.$ (9)
We choose $E_{0}=-N\omega_{2}-N(N-1)g_{22}/2$ so that $\hat{H}\ket{0,N}=0$ for convenience. The choice of $E_{0}$ will not affect the net output of any cyclic process, for example the performance of an engine cycle. The detuning $\Delta$ for two spatial modes can be controlled via tuning the trapping potential [47], and for two spin states via tuning of a Zeeman field [48, 49]. As will be discussed below, we require thermal states that have a Gaussian (or approximately Gaussian) distribution over states |$\ket{n,N-n}$, which requires $g>0$. Hence the interaction strengths must satisfy $g_{11}+g_{22}>2g_{12}$, i.e. the average intramode interaction strength $(g_{11}+g_{22})/2$ must exceed the intermode interaction strength $g_{12}$. For two spatially separated modes we have $g_{12}=0$ and therefore $g>0$ is ensured as long as the intramode interactions $g_{ii}$ are positive, as realised for example in 87Rb and 23Na condensates [50, 51].
@@
the partition function. Due to particle interactions, the boson number distribution of a thermal state is approximately Gaussian (when $\beta g>0$ and $-\beta\Delta\gg\sqrt{\beta g}$), in contrast to the Planck distribution of non-interacting bosons. This feature will be essential for the efficacy of our protocol as it will enable close matching between the initial and thermal boson number distributions. We assume $\beta g\ll 1$; noting that $gN\sim\mu$, where $\mu$ is the BEC chemical potential, this is ensured when $\beta\mu/N\ll 1$, a condition satisfied in most realisations of dilute gaseous BECs [52, 53]. We can then approximate the sum in Eq. (12) by an integral. We also assume $N\gg 1$ and $\langle\hat{a}^{\dagger}\hat{a}\rangle\ll N$, i.e. most of the particles are in the $\hat{b}$ mode throughout the protocol. We can then set $N\rightarrow\infty$ and approximate |$\ket{n,N-n}$ by a harmonic oscillator number state, which simplifies our calculations [54, 55]. Evaluating Eq. (12) with these two approximations gives
$Z\approx\sqrt{\frac{\pi}{\beta g}}e^{\beta(g-\Delta)^{2}/(4g)}\left(\frac{1+\operatorname{erf}\left(\frac{\beta(g-\Delta)}{2\sqrt{\beta g}}\right)}{2}\right).$ (13)
@@ III Results
$p_{n}^{G}=\frac{|\alpha|^{2n}e^{-|\alpha|^{2}}}{n!}\approx\frac{1}{\sqrt{2\pi|\alpha|^{2}}}e^{-(n-|\alpha|^{2})^{2}/(2|\alpha|^{2})}.$ (15)
@@
For more general parameters, the work output will have both a contribution from coherence and a “classical” contribution. Below we present an explicit protocol to show how work can be extracted from the coherence of a GCS. We then show how the work output depends on particle number and temperature, identifying a regime where the work output is dominated by the contribution from coherence. Next we show how the work output can be improved by squeezing the initial state and finally we discuss the role of entanglement in the system.
@@ III.1 Protocol to extract work from a GCS
where $S_{i}^{V}(S_{f}^{V})$ are the initial (final) von Neumann entropies of the system.
For the protocol in Fig. 1 we have $S^{V}_{i}=0$ since the initial state is pure and $S^{V}_{f}=S^{E}_{f}$ since the final state is thermal. Here $S^{E}_{i}(S^{E}_{f})$ are the energy entropies for the initial (final) state. The total entropic change $S^{V}_{f}-S^{V}_{i}$ can now be decomposed into a classical contribution $S^{E}_{f}-S^{E}_{i}$ and a quantum contribution $S^{E}_{i}-S^{V}_{i}=D(\rho||\rho_{\mathrm{ed}})$ arising from coherence, Eq. (3). The work then consists of two parts [11, 75, 13], a quantum part as defined in Eq. (4), as well as a classical part:
$\begin{split}W_{\mathrm{quant}}=&\beta^{-1}(S^{E}_{i}-S^{V}_{i})\\ W_{\mathrm{class}}=&U_{i}-U_{f}-\beta^{-1}(S^{E}_{i}-S^{E}_{f}).\end{split}$ (18)
Note that a projection of the initial state Eq. (14) onto the energy eigenbasis prior to work extraction, i.e. $\rho\to\rho_{\mathrm{ed}}$, would result in $W_{\mathrm{quant}}=0$, but would not change $W_{\mathrm{class}}$. The protocol in Fig. 1 avoids such projection, allowing $W_{\mathrm{quant}}$ to be extracted.
@@ III.2 Classical and quantum contributions to the work extracted from a GCS
With an initial state (14), the outputs $W_{\mathrm{quant}}$ and $W_{\mathrm{class}}$ in Eq. (18) can be calculated analytically. We choose $|\alpha|$ so that the GCS has mean boson number equal to the thermal state,
$|\alpha|^{2}=\langle\hat{n}\rangle_{\mathrm{therm}}\equiv\langle\hat{n}\rangle.$ (19)
Here $\hat{n}=\hat{a}^{\dagger}\hat{a}$ and a “therm” subscript denotes expectation values for state $\rho_{\mathrm{therm}}$, Eq. (10). We have checked that this choice of $|\alpha|$ minimises $W_{\mathrm{class}}$ to a very good approximation over the parameters explored. The energetic and entropic terms in Eq. (18) are
$\begin{split}U_{i}-U_{f}=&g\left(\langle\hat{n}\rangle-\langle\delta\hat{n}^{2}\rangle_{\mathrm{therm}}\right),\\ S^{E}_{i}=&\frac{1}{2}+\frac{1}{2}\ln\left(2\pi\langle\hat{n}\rangle\right)+O\left(\frac{1}{\langle\hat{n}\rangle}\right),\\ S^{E}_{f}=&\frac{1}{2}+\frac{1}{2}\ln\left(2\pi\langle\delta\hat{n}^{2}\rangle_{\mathrm{therm}}\right)+S_{\mathrm{NG}},\end{split}$ (20)
@@
with $u=\frac{\beta(g-\Delta_{0})}{2\sqrt{\beta g}}$ and $G(u)=\frac{e^{-u^{2}}}{\sqrt{\pi}(1+\operatorname{erf}u)}=O(e^{-\beta\Delta_{0}^{2}/(4g)})$. The entropic correction $S_{\mathrm{NG}}$ arises since the thermal distribution is not a perfect Gaussian, and is only important for $u\lesssim 1$.
@@
Again, the $G(u)$ terms are non-Gaussian corrections, which are only important for $u\lesssim 1$. Note for $u\gg 1$ we have $u^{2}\approx\langle\hat{n}\rangle_{\mathrm{therm}}^{2}/(2\langle\delta\hat{n}^{2}\rangle_{\mathrm{therm}})$ and hence $u^{-1}$ is essentially the normalized second-order correlation function of the thermal bosonic field [76]. The thermal mean, variance and third central moment $\langle\delta\hat{n}^{3}\rangle=\langle(\hat{n}-\langle\hat{n}\rangle)^{3}\rangle$ are shown in the inset to Fig. 2(a). We focus on cases $\langle\hat{n}\rangle_{\mathrm{therm}}\gtrsim 1$, which is ensured when the initial detuning is less than the reservoir temperature,
@@
Figure 2: (a) Work extracted from a two-mode BEC undergoing the protocol in Fig. 1 for different mean bosonic numbers $\langle\hat{n}\rangle$. Work extracted for an initial GCS (blue curves) divided into quantum $\beta W_{\mathrm{quant}}$ (blue solid curve) and classical $\beta W_{\mathrm{class}}$ (blue dotted curve) contributions, see Eq. (18). The work output is dominated by $W_{\mathrm{quant}}$ for $\langle\hat{n}\rangle\lesssim n_{G}$, with $n_{G}$ given by Eq. (27) and marked by a vertical dashed line. The number state distributions for the GCS and thermal states coincide almost perfectly when $\langle\hat{n}\rangle=1/(2\beta g)$, at which point $W_{\mathrm{class}}\approx 0$ is minimised (dip in blue dotted curve). Inset: First three cumulants of the thermal boson number distribution, with $\langle\delta\hat{n}^{k}\rangle=\langle(\hat{n}-\langle\hat{n}\rangle)^{k}\rangle$ for $k=2,3$. The distribution is Gaussian for $u\gg 1$, in which case $\langle\hat{n}\rangle_{\mathrm{therm}}\approx u/\sqrt{\beta g}$, $\langle\delta\hat{n}^{2}\rangle_{\mathrm{therm}}\approx 1/(2\beta g)$ and higher order cumulants are zero. (b) (i)-(iv) Number state distributions for different mean bosonic numbers (four panel values are marked by crosses in (a)) for a GCS (blue curve) and a thermal state (black dashed curve). The GCS distribution is narrower (wider) than the thermal distribution for $\langle\hat{n}\rangle$ smaller (larger) than $\langle\delta\hat{n}^{2}\rangle_{\mathrm{therm}}\approx 1/(2\beta g)\approx 501$, which results in non-zero $W_{\mathrm{class}}$. Pale red curves in (a) and (b): as for the GCS but for a squeezed state. These are included here for comparison and will be discussed in Sec. III.3. The GCS and squeezed distributions coincide in (ii). All results are for $\beta g=10^{-3}$. Different $\langle\hat{n}\rangle$ correspond to different $\beta\Delta_{0}$, which in (b) are: (i) $-\beta\Delta_{0}=0$, (ii) $-\beta\Delta_{0}=1$, (iii) $-\beta\Delta_{0}=50$, (iv) $-\beta\Delta_{0}=10^{3}$.
Analytic expressions for $W_{\mathrm{class}}$ and $W_{\mathrm{quant}}$ follow from Eq. (18)–(23) and are plotted in Fig. 2(a) as a function of $\langle\hat{n}\rangle$. Note $\langle\hat{n}\rangle$ is itself a function of $\Delta_{0}$, according to Eq. (22), and hence varying $\langle\hat{n}\rangle$ corresponds to an implicit variation of $\Delta_{0}$. The minimum of the classical work occurs when $p_{n}^{G}$ and $p_{n}^{\mathrm{therm}}$ are as close as possible, see Fig. 2(b). For large $\langle\hat{n}\rangle$ both distributions are approximately Gaussian and hence will be equal when their respective means and variances are equal. Since $p_{n}^{G}$ has equal mean and variance, this requires
@@
This occurs when $\beta\Delta_{0}\approx-1$, at which point $\langle\hat{n}\rangle_{\mathrm{therm}}=\langle\delta\hat{n}^{2}\rangle_{\mathrm{therm}}\approx 1/(2\beta g)$. At this optimal point, $W_{\mathrm{class}}=O(\langle\hat{n}\rangle^{-1})$ and the work output almost exclusively consists of $W_{\mathrm{quant}}$. (Note our choice of units: plotting work in units of $\beta^{-1}$ is equivalent to plotting the information lost if the work was dissipated as heat, according to Landauer’s principle [77].)
@@
$\begin{split}W_{\mathrm{class}}&\approx g\langle\hat{n}\rangle-\frac{1}{2\beta}-\frac{1}{2\beta}\ln\left(2\beta g\langle\hat{n}\rangle\right),\\ W_{\mathrm{quant}}&\approx\frac{1}{2\beta}+\frac{1}{2\beta}\ln\left(2\pi\langle\hat{n}\rangle\right).\end{split}\hskip 28.45274pt(u\gg 1)$ (26)
@@
see Fig. 2(a). Here $\Omega_{k}$ is the $k$th branch of the omega function  [^1] Also known as the Lambert $W$-function, denoted by $W_{k}$. We avoid this notation to avoid confusion with work output.. The function $-\Omega_{-1}(-x)$ satisfies $1\leq-\Omega_{-1}(-x)<\infty$ for $x\in(0,\exp(-1)]$ and monotonically decreases with increasing $x$, with $-\Omega_{-1}(-x)\sim\ln\left(\frac{1}{x}\ln\frac{1}{x}\right)$ for small $x$ [80]. Hence, ignoring logarithmic corrections, $n_{G}$ decreases with increasing $\beta g$ as $n_{G}\sim 1/(\beta g)$. For $\langle\hat{n}\rangle>n_{G}$, the classical work output grows extensively due to the extensive growth of $U_{i}-U_{f}$, see Eq. (26). The quantum work also increases with $\langle\hat{n}\rangle$ due to increased coherence in the initial state. The increase, however, is subextensive, with $W_{\mathrm{quant}}\sim\frac{1}{2\beta}\ln\langle\hat{n}\rangle$. For $\langle\hat{n}\rangle<\langle\delta\hat{n}^{2}\rangle_{\mathrm{therm}}$ (thermal bunching), the number fluctuations of the initial state are less than the final state, see Fig. 2(b). Although this increases $W_{\mathrm{class}}$, we still have $W_{\mathrm{quant}}>W_{\mathrm{class}}$ as long as $\langle\hat{n}\rangle\gg 1$, see Fig. 2(a).